OpenStudy (anonymous):

Find an equation for the line: through (–4, 6) and parallel to y = -3x + 4

OpenStudy (anonymous):

NOTE: The equation of 2 parallel lines will always have the same slope. Given y = -3x+4, this is written in slope intercept form (i.e y=mx+b, where m=slope, b= y-intercept). 1. From the equation: y = -3x+4: slope(m)=-3. The slope (m) of the parallel line is also m=-3. 2. Use the point-slope form: \[y-y _{1}=m(x-x _{1})\] Slope (m)=-3 and given the point: (-4,6) from the point \[x _{1}=-4, y _{1}=6\] equation: y-6= -3(x-(-4)) y-6 = -3(x+4) y-6 = -3x-12 (solve for y by adding 6 to both sides) y = -3x-6 (answer in slope-intercept form) To write answer in Standard form (ax+by+c=0), where a is positive and all the coefficients are integers (not fractions or decimals) we have y-6 = -3x-12 (move x term to the left side: by adding 3x to both sides) 3x+y-6 = -12 (add 12 to both sides) 3x+y+6 = 0, in standard form

OpenStudy (anonymous):

Any questions about this?

OpenStudy (anonymous):

No, haha. Thanks!